# Continual Learning of Large Language Models: A Comprehensive Survey

### v2

@@ 5.2. Datasets
6. Discussion
In this section, we delve into the intersection of conventional computational patterns in continual learning and the training and deployment of large language models (LLMs). We begin by examining intriguing properties that arise during continual learning with LLMs. Next, we explore the evolving roles of three types of incremental learning within the context of LLMs. Following this, we contrast the roles of memory in continual LLMs with those in traditional continual learning. Finally, we conclude with a concise overview of promising directions for future research in this area.
6.1. Intriguing Properties Emergent in Continual LLMs
Beyond the well-established resilience of pre-trained large language models (LLMs) against catastrophic forgetting compared to downstream-specific models (ke2021achieve, ; tao2022can, ; luo2023investigating, ; zheng2023learn, ; mehta2023empirical, ), there is a notable lack of exploration into other intriguing properties of LLMs when trained continually. While investigations into the emergent capabilities of continuously trained LLMs have attracted attentions from the community to a certain extent, they remain relatively limited. For instance, in (yang2024reawakening, ), it is observed that when fine-tuned sequentially and cyclically on a series of documents, large models exhibit a phenomenon known as “anticipatory recovering”. This refers to the LLMs’ ability to recover forgotten information on documents even before encountering them again. This suggests that LLMs may possess the capability of sequential memorization, which could pave the way for research into memory replay and more complex structured learning environments as model parameters scale up.
6.2. Conventional Types of Incremental Learning
As mentioned in Section 2.2, three types of incremental learning are prevalent (van2022three, ). Among them, class-incremental learning (CIL) has historically attracted significant attention from the community (rebuffi2017icarl, ; wu2019large, ). However, in the context of continually pre-training and adapting large language models (LLMs), we observe a decreased interest in CIL but an increased focus on task-incremental learning (TIL) and domain-incremental learning (DIL). Given that language models are inherently designed for content generation and are pre-trained with the pretext generative task of next-word prediction, it is natural to emphasize the patterns of generative tasks and integrate the traditional CIL paradigm into the broader framework of language modeling, discarding the incremental classification head (shao2023class, ; dalessandro2023multimodal, ; cao2024generative, ). For instance, in Vocabulary-Aware Label Generation (VAG), CIL is redefined as the task of continual label generation. This approach utilizes a pre-trained encoder-decoder language model to generate class labels (shao2023class, ). Meanwhile, in the Generative Multi-modal Model (GMM) for CIL (cao2024generative, ), image patches and prompts are concatenated and fed into the language model to generate classification results.
However, the declining attention to the conventional CIL paradigm does not suggest that these techniques are not impactful in the field of continual learning for LLMs. Nonetheless, many current research endeavors unwittingly employ such techniques, indicating their widespread adoption in various applications. For example, techniques such as vocabulary expansion (amba2021dynamic, ; cossu2022continual, ) can be seen as an extension of expanding the classification head in CIL. These CIL techniques can be further integrated into systems like Lifelong-MoE (chen2023lifelong, ), where adding a new expert to the transformer blocks requires updating the gating function to include the routing of the newly added expert. The EProj framework, described in (he2023continual, ), employs a similar architecture, incorporating linear projectors for new domains alongside a selector module trained to route these projectors. Since the aforementioned sub-modules operate on the principles of CIL, previously validated techniques can be directly applied.
The importance of domain-incremental learning (DIL) is self-evident, given the shared task definition and input-output format in continual pre-training (CPT) and domain-adaptive pre-training (DAP). As dynamically expanding token vocabularies can pose additional challenges, it is natural to focus on understanding distributional shifts within the input corpus while keeping the vocabulary fixed. On another front, task-incremental learning (TIL) attracts significant interest due to its potential for personalizing LLM services. For instance, users may desire options for selecting domain-specific experts, thereby making task IDs available throughout inference time (huang2023lorahub, ; wistuba2023continual, ). Additionally, TIL plays a crucial role in instruction tuning, where instructions can be seen as natural-language-encoded task information (scialom2022fine, ; huang2024mitigating, ; mok2023large, ; he2024dont, ; yin2022contintin, ; wang2023orthogonal, ; zhao2024sapt, ; wang2024inscl, ). It is worth noting that the boundary between TIL and DIL becomes somewhat blurred in continual instruction tuning. Language models demonstrate the capability to infer domain information for unseen instructions, suggesting a convergence of TIL and DIL in certain contexts.
6.3. Roles of Memory in Continual LLMs
Previous continual learning research, drawing inspiration from human learning patterns, primarily emphasizes the storage efficiency of past data. The setting of continual learning with limited memory size has garnered significant attention from the community. However, this focus may no longer hold true in the context of continual LLMs. In the direction of relaxing memory constraints, institutions with access to training data may opt to retain full access without restricting memory size, given that the cost of memory storage is more than affordable. In such scenarios, as highlighted in (verwimp2024continual, ), the challenge shifts from storage efficiency to computational efficiency. To achieve continual learning goals, models must efficiently adapt to new data (efficient adaptation) and select key experiences for replay (efficient replay) (xie2023efficient, ; jin2024model, ). Therefore, it is essential to reassess the existing memory constraint and prioritize optimizing computational efficiency for continual learning of LLMs by restricting the number of updates and the number of FLOPs (prabhu2023computationally, ; wang2022sparcl, ).
On the other end of the spectrum, studies with tightened memory constraints remain vital in modern continual learning of LLMs. As shown in Fig. 1, upstream suppliers of LLMs typically do not provide training data with the released model weights. Consequently, consumers must adapt these models to downstream data without access to the actual replay data. Various rehearsal-free continual strategies are applied in this scenario, such as collecting data examples from alternate sources (rozière2024code, ; colombo2024saullm7b, ; wu2023pmc, ; Azerbayev2023LLEMMA, ), leveraging the generative capabilities of LLMs to produce pseudo-examples for replay (qin2021lfpt5, ), and implementing regularization techniques in the parameter space (ke2022continual-pre, ; rongali2021continual, ). Continual learning under the strict memory constraint is also driven by data privacy concerns, where preserving data on the server side is prohibited. In these scenarios, researchers must rely on online continual learning methods (cai2021online, ; mai2022online, ; prabhu2023online, ), where data examples are only utilized for training as they arrive in a stream, and numerous efforts are already underway to develop LLMs capable of operating under these constraints (yang2022continual, ; wang2022online, ; bornschein2024transformers, ).
6.4. Prospective Directions
Efficient Replay for Knowledge Retention for Continual LLMs. Computational resources for training large-scale LLMs are often limited. While the storage budget can theoretically be infinite (Section 6.3), replaying past experiences without specific design can lead to inefficient updates in current domain learning, resulting in slow convergence. Beyond sparse replay solutions that control data mixture ratios (lin2023geogalactica, ; rozière2024code, ; Yang2023PLLaMa, ), there is ongoing exploration of efficient replay for continual LLMs. For example, KPIG (he2024dont, ) enhances replay efficiency by calculating Key-part Information Gain (KPIG) on masked segments, enabling the dynamic selection of replay data. In (jin2024model, ), a pioneering effort introduces a forgetting forecasting mechanism based on output changes during adaptation, later used for selective replay in continual model refinement (CMR). It has been verified in this work that filtering replay samples based on their tendency to forget significantly improves knowledge retention rates for continual LLMs. However, more sophisticated and accurate data mixing strategies and efficient replay sample selection mechanisms are much needed, e.g., a dynamic data mixing ratio throughout the training process. Hence we mark this practical direction of efficient replay for LLMs a significant research focus in the future.
Continual LLMs with Controllable Memory. The long-term memory inherent in the whole set of parameters of LLMs often lacks interpretability and explicit manipulability, which is crucial in certain application areas. For instance, consider a scenario where a supplier collects data from customers under their consent and continually utilizes this data to update LLMs. However, if some users later revoke their consent, the knowledge acquired by the trained model from that portion of data must also be revoked. With a continually pre-trained large-scale LLM, the only solution is to roll back to a previous model version predating the inclusion of these users’ data and retrain the model from that point onward. This example of “machine unlearning” (bourtoule2020machine, ; nguyen2022survey, ) vividly illustrates the benefits of equipping LLMs with an external, controllable memory. As part of continual model refinement (CMR), memory systems for continual learning have been explored in several studies. Larimar (das2024larimar, ) suggests integrating the Kanerva Machine (wu2018kanerva, ) as an episodic memory for multi-fact model editing. This memory system supports basic operations like writing, reading, and generating, as well as advanced operations such as sequential writing and forgetting. It enables one-shot knowledge updates without costly retraining or fine-tuning. Additionally, other memory systems like Hopfield Networks (ramsauer2021hopfield, ; pourcel2022online, ) hold promise for future investigation.
Continual LLMs with Custom Preferences. Customizing user preferences is critical for LLMs, especially in service-oriented contexts. Users often require different trade-offs between domain expertise, ethics, values, or tones of expression. Efficiently building customized LLMs for individual users and offering flexible adjustment options is a challenging task. Early attempts in this direction include Imprecise Bayesian Continual Learning (IBCL), which, under certain assumptions, guarantees the generation of Pareto-optimal models based on user preferences by combining two model posteriors in the parameter space (lu2023ibcl, ). While empirical validation is limited in scale, this approach paves the way for future research in this area.
7. Conclusion

### v3

@@ 5. Evaluation Protocols and Datasets
6. Discussion
6.1. Intriguing Properties Emergent in Continual LLMs
Beyond the well-established resilience of pre-trained large language models (LLMs) against catastrophic forgetting compared to downstream-specific models (ke2021achieve, ; tao2022can, ; luo2023investigating, ; zheng2023learn, ; mehta2023empirical, ), there is a notable lack of exploration into other intriguing properties of LLMs when trained continually. In (yang2024reawakening, ), it is observed that when fine-tuned sequentially and cyclically on a series of documents, large models exhibit a phenomenon known as “anticipatory recovering.” This refers to the LLMs’ ability to recover forgotten information on documents even before encountering them again. This suggests that LLMs may possess the capability of sequential memorization, which could pave the way for research into more complex structured learning environments as model parameters scale up.
6.2. Conventional Types of Incremental Learning
As mentioned in Section 2.2.1, three types of incremental learning are prevalent (van2022three, ). Among them, class-incremental learning (CIL) has historically attracted significant attention from the community (rebuffi2017icarl, ; wu2019large, ). However, in the context of continually pre-training and adapting large language models (LLMs), we observe a decreased interest in CIL but an increased focus on task-incremental learning (TIL) and domain-incremental learning (DIL). Given that language models are inherently designed for content generation and are pre-trained with the pretext generative task of next-word prediction, it is natural to emphasize the patterns of generative tasks and integrate the traditional CIL paradigm into the broader framework of language modeling, discarding the incremental classification head (shao2023class, ; cao2024generative, ). However, the declining attention to CIL does not suggest that it is not impactful in the field of continual learning for LLMs. Techniques such as vocabulary expansion (amba2021dynamic, ; cossu2022continual, ) and learning routing function in the MoE system (chen2023lifelong, ) can be seen as an extension of expanding the classification head in CIL, and previously validated techniques of CIL can be directly applied.
The importance of DIL is self-evident, given the shared task definition and input-output format in continual pre-training (CPT) and domain-adaptive pre-training (DAP). On the other hand, TIL attracts significant interest as it plays a crucial role in instruction tuning, where instructions can be seen as natural-language-encoded task indices (scialom2022fine, ; huang2024mitigating, ; mok2023large, ; he2024dont, ; yin2022contintin, ; wang2023orthogonal, ; zhao2024sapt, ; wang2024inscl, ). It is worth noting that the boundary between TIL and DIL becomes somewhat blurred in continual instruction tuning. Language models demonstrate the capability to infer domain information for unseen instructions, suggesting a convergence of TIL and DIL in certain contexts.
6.3. Roles of Memory in Continual LLMs
Previous continual learning research, drawing inspiration from human learning patterns, primarily emphasizes the storage efficiency of past data. However, this focus may no longer hold true in the context of continual LLMs. In the direction of relaxing memory constraints, institutions with access to training data may opt to retain full access without restricting memory size, given that the cost of memory storage is more than affordable. In such scenarios, as highlighted in (verwimp2024continual, ), the challenge shifts from storage efficiency to computational efficiency. To achieve continual learning goals, models must efficiently adapt to new data (efficient adaptation) and select key experiences for replay (efficient replay) (xie2023efficient, ; jin2024model, ). Therefore, it is essential to reassess the existing memory constraint and prioritize optimizing computational efficiency for continual learning of LLMs by restricting the number of updates and FLOPs (prabhu2023computationally, ; wang2022sparcl, ).
On the other end of the spectrum, studies with tightened memory constraints remain vital in modern continual learning of LLMs. As shown in Fig. 1, upstream suppliers of LLMs typically do not provide training data with the released model weights. Consequently, consumers must adapt these models to downstream data without access to the actual replay data. Various rehearsal-free continual strategies are applied in this scenario, such as collecting data examples from alternate sources (rozière2024code, ; colombo2024saullm7b, ; wu2023pmc, ; Azerbayev2023LLEMMA, ), leveraging the generative capabilities of LLMs to produce pseudo-examples for replay (qin2021lfpt5, ), and implementing regularization techniques in the parameter space (ke2022continual-pre, ; rongali2021continual, ). Continual learning under the strict memory constraint is also driven by data privacy concerns, where preserving data on the server side is prohibited. In these scenarios, researchers must rely on online continual learning methods (mai2022online, ; prabhu2023online, ), where data examples are only utilized for training as they arrive in a stream, and numerous efforts are already underway to develop LLMs capable of operating under these constraints (bornschein2024transformers, ).
6.4. Prospective Directions
Efficient Replay for Knowledge Retention for Continual LLMs. While the storage budget can theoretically be infinite (Section 6.3), replaying past experiences without specific design can lead to inefficient updates in current domain learning, resulting in slow convergence. Beyond sparse replay solutions that control data mixture ratios (lin2023geogalactica, ; rozière2024code, ; Yang2023PLLaMa, ), there is ongoing exploration of efficient replay for continual LLMs. For example, KPIG (he2024dont, ) enhances replay efficiency by calculating Key-Part Information Gain (KPIG) on masked segments, enabling the dynamic selection of replay data. (jin2024model, ) introduces a forgetting forecasting mechanism based on output changes during adaptation, later used for selective replay in continual model refinement (CMR). More sophisticated and accurate data mixing strategies and efficient replay sample selection mechanisms are needed and hence we mark it as a significant research focus in the future.
Continual LLMs with Controllable Memory. The long-term memory inherent in the whole set of parameters of LLMs often lacks interpretability and explicit manipulability, which is crucial in certain application areas such as machine unlearning (bourtoule2020machine, ), where the continually pre-trained models need to constantly roll back to a previous version predating the inclusion of the revoked data and retrain the model from that point onward. This example illustrates the benefits of equipping LLMs with an external, controllable memory. As part of continual model refinement (CMR), memory systems for continual learning have been explored in several studies. Larimar (das2024larimar, ) suggests integrating the Kanerva Machine (wu2018kanerva, ) as an episodic memory for multi-fact model editing. This memory system supports basic operations like writing, reading, and generating, as well as advanced operations such as sequential writing and forgetting. It enables one-shot knowledge updates without costly retraining or fine-tuning. Other memory systems like Hopfield Networks (ramsauer2021hopfield, ) hold promise for future investigation as well.
Continual LLMs with Custom Preferences. In service-oriented contexts, users often require different trade-offs between domain expertise, ethics, values, or tones of expression. Efficiently building customized LLMs for individual users and offering flexible adjustment options is a challenging task. Early attempts in this direction include Imprecise Bayesian Continual Learning (IBCL), which, under certain assumptions, guarantees the generation of Pareto-optimal models based on user preferences by combining two model posteriors in the parameter space (lu2023ibcl, ). While empirical validation is limited in scale, this approach paves the way for future research in this area.
7. Conclusion
Supplementary Material
Appendix A Preliminaries
In this section, we provide an overview of the fundamental concepts of large language models (LLMs) and continual learning (CL) We begin by introducing the notation used in this paper. Subsequently, we discuss the pre-training and downstream adaptation of LLMs, as well as mainstream LLM families (Appendix A.1), followed by an introduction to basic continual learning techniques studied by the community (Appendix A.2).
A.1. Large Language Models
A.1.1. Pre-Training of LLMs
A.1.2. Adaptation of LLMs
LLMs are primarily trained to generate linguistically coherent text. However, this training may not align with human values, preferences, or practical needs. Furthermore, the pre-training data can be outdated, leading to knowledge cutoffs or inaccuracies. To address these issues, various computational paradigms such as Instruction Tuning (IT) (zhang2024instruction, ), Model Refinement (MR) (de2021editing, ), and Model Alignment (MA) (ouyang2022rlhf, ; rafailov2024dpo, ) have been proposed. These approaches adapt LLMs to better meet diverse downstream tasks and user requirements.
Definition A.0 (Instruction Tuning, IT).
Remark 0.
Definition A.0 (Model Refinement, MR).
A.2. Continual Learning
Humans gradually accumulate knowledge and skills across tasks without significant performance decline on previous tasks (mcclelland1995there, ; kandel2000principles, ; pallier2003brain, ; mccaffary2021towards, ). In contrast, machine learning models are usually data-centric, minimizing the training loss on the subsequent tasks will cause the model fail on the old ones, which phenomenon is phrased as “catastrophic forgetting”. Addressing this challenge is a focal point in continual learning research. The problem of efficiently adapting models to a sequence of tasks without forgetting is extensively studied in the continual learning community (pentina2016theoretical, ; chen2018lifelong, ; van2022three, ; wang2024comprehensive, ). These studies are typically conducted under the following memory constraint of CL.
Definition A.0 (Memory Constraint of Continual Learning).
Remark 0.
In early stages of CL, works mostly focused on the strong memory constraint (kirkpatrick2017overcoming, ; li2017learning, ; aljundi2018memory, ; lomonaco2020rehearsalfree, ); as the research field progresses, more focus was put on relaxing the memory constraint to a small buffer for replay (rebuffi2017icarl, ; chaudhry2019tiny, ; buzzega2020dark, ; shi2024unified, ); some modern CL works completely discard the memory constraint but put focus on the computational budget (prabhu2023online, ; verwimp2024continual, ).
A.2.1. Three Types of Continual Learning
Definition A.0 (Task-Incremental Learning, TIL).
Suppose $T$ task distributions $\{{\mathcal{T}}_{t}\}_{t=1}^{T}$ come in as a sequence, where ${\mathcal{T}}_{t}$ denotes the joint distribution over the $t$-th task’s input space and the label space $({\mathcal{X}}_{t},{\mathcal{Y}}_{t})$. Denote ${\mathcal{X}}\triangleq\bigcup_{t=1}^{T}{\mathcal{X}}_{t}$ and ${\mathcal{Y}}\triangleq\bigcup_{t=1}^{T}{\mathcal{Y}}_{t}$ as the union of the input and label spaces, respectively. Under the memory constraint defined in Definition A.3, Task-Incremental Learning (TIL) aims to find the optimal hypothesis $h^{*}:{\mathcal{X}}\times[T]\rightarrow{\mathcal{Y}}$ that satisfies:
Definition A.0 (Domain-Incremental Learning, DIL).
Suppose $T$ domain distributions $\{{\mathcal{D}}_{t}\}_{t=1}^{T}$ come in as a sequence, where ${\mathcal{D}}_{t}$ denotes the $t$-th joint distribution over the shared input space and label space $({\mathcal{X}},{\mathcal{Y}})$. Under the memory constraint defined in Definition A.3, Domain-Incremental Learning (DIL) aims to find the optimal hypothesis $h^{*}:{\mathcal{X}}\rightarrow{\mathcal{Y}}$ that satisfies:
Definition A.0 (Class-Incremental Learning, CIL).
Suppose $T$ task distributions $\{{\mathcal{T}}_{t}\}_{t=1}^{T}$ come in as a sequence, where ${\mathcal{T}}_{t}$ denotes the joint distribution over the $t$-th task’s input space and the label space $({\mathcal{X}}_{t},{\mathcal{Y}}_{t})$. Denote ${\mathcal{X}}\triangleq\bigcup_{t=1}^{T}{\mathcal{X}}_{t}$ and ${\mathcal{Y}}\triangleq\bigcup_{t=1}^{T}{\mathcal{Y}}_{t}$ as the union of the input and label spaces, respectively. Under the memory constraint defined in Definition A.3, Class-Incremental Learning (CIL) aims to find the optimal hypothesis $h^{*}:{\mathcal{X}}\rightarrow[T]\times{\mathcal{Y}}$ that satisfies:
Remark 0.
A.2.2. Techniques of Continual Learning
The objectives for past domains are often challenging to measure or optimize due to the memory constraints (Definition A.3). Therefore, the core of designing CL algorithms lies in identifying a proxy learning objective for the first term without violating the memory constraint. Existing CL techniques can be roughly categorized into 5 groups: (i) replay-based, (ii) regularization-based, (iii) architecture-based, (iv) optimization-based, and (v) representation-based (wang2024comprehensive, ). Here, we provide a concise yet comprehensive introduction to the first three categories of continual learning techniques, as they find extensive application in continually learning large language models.
where ^$\widehat{{\mathcal{L}}}_{S}$ denotes the empirical loss term evaluated on the set of examples $S$. Often regarded as a simplistic solution to CL, replay-based methods may theoretically lead to loose generalization bounds (shi2024unified, ). Despite this, they are valued for their simplicity, stability, and high performance, even with a small episodic memory (chaudhry2019tiny, ; riemer2018learning, ). For instance, DER++ (buzzega2020dark, ) demonstrates consistent performance enhancement by replaying a small set of past examples along with their logits (known as dark experience replay). ESM-ER (sarfraz2023error, ) introduces error sensitivity modulation (ESM) to mitigate abrupt representational drift caused by high-error new examples. A significant focus in replay-based CL is enhancing sample efficiency for buffer maintenance. For instance,  (rebuffi2017icarl, ) prioritizes exemplar selection based on herding to accurately model class mean throughout class-incremental learning. (zhao2022memory, ) propose storing low-fidelity examples to achieve memory-efficient exemplar set maintenance.
Architecture-Based Methods. Expanding the network architecture dynamically to assimilate new knowledge is deemed the most efficient form of CL (wang2022learning, ; wang2022dualprompt, ). This method primarily tackles adaptation challenges and can achieve zero-forgetting when task IDs are available during inference or can be correctly inferred (gururangan2022demix, ; wistuba2023, ). However, due to the difficulty of task ID inference, architecture expansion is predominantly utilized in TIL but is scarcely explored in DIL or CIL. Progressive Neural Networks (PNN) (rusu2016progressive, ) proposes learning laterally connected neurons as new tasks arise, ensuring non-forgetting and enabling transfer of previously learned neurons for future tasks. In conjunction with pre-trained backbone large models like ViT (dosovitskiy2020image, ), CoLoR (wistuba2023, ) trains various low-rank adaptation (LoRA) (hu2021lora, ) modules for different tasks. It estimates and stores prototypes for each task and utilizes the natural clustering ability of the pre-trained model during testing to infer task IDs, selecting the corresponding LoRA component for prediction generation. In the domain of continual LLMs, architecture expansion has resurged in popularity following the rise of parameter-efficient fine-tuning (PEFT) (shazeer2017outrageously, ; hu2021lora, ; dettmers2023qlora, ), a topic we will delve into shortly (yang2024moral, ; wang2023orthogonal, ; li2024examining, ; jang2022towards, ; jin2022lifelong, ; paul2024ircoder, ; yan2023af, ; wu2024llama, ).
Appendix B Evaluation Protocols and Datasets
In Appendix B.1, we review common continual learning evaluation metrics and provide formal definitions. In Appendix B.2, we introduce metrics designed specifically for continual LLMs. Finally, in Appendix B.3, we outline the datasets available for each discussed topic.
B.1. Evaluation Metrics of Continual Learning
(11) $\operatorname{OP}_{t}$ $\triangleq\tfrac{1}{t}\sum_{i=1}^{t}P_{t,i}.$
As noted in (shi2024unified, ), the OP corresponds to the primary optimization objective defined in Definition A.4, A.5, and A.6. In much of the continual learning literature, once all $T$ tasks are completed, the final $\operatorname{OP}$ ($\operatorname{OP}_{T}$) is reported, with the subscript T often omitted for brevity. In some works, OP is weighted by the importance of tasks $\widetilde{\operatorname{OP}}\triangleq\tfrac{1}{T}\sum_{i=1}^{T}w_{i}P_{t,i}$, where $w_{i}=N_{i}/\sum_{j=1}^{T}N_{j}$ represents the ratio of data. In some literature, $\widetilde{\operatorname{OP}}$ is referred to as “example accuracy” (chen2024parameterizing, ), “whole accuracy” (song2023conpet, ), or “edit success rate” in CMR (hartvigsen2023aging, ).
Typically, researchers report the average forgetting $F=F_{T}$ at the end of the entire training process. Forgetting quantifies the impact of learning new tasks on previously acquired knowledge. Ideally, a robust continual learning framework should achieve Backward Transfer (BWT), where learning new tasks enhances performance on prior tasks. This enhancement is typically measured by negating the forgetting, thus indicating an improvement in performance on earlier tasks. The concepts of Forgetting and Backward Transfer underpin various evaluation metrics, such as knowledge retention (jin2022lifelong, ), performance on unchanged knowledge (jang2022temporalwiki, ), average increased perplexity (AP+(qin2022elle, ), and test and edit retention rate in CMR (hartvigsen2023aging, ).
where $b_{i}$ is the baseline performance of the model evaluated on task $i$ before undergoing continual learning. Strictly speaking, the definition of $b_{i}$ is not the same as defined in the previous work (lopez2017gradient, ; shi2024unified, ), where it is used to denote the performance of a random initialization of the model. Additionally, we extend the notation of forward transfer in the vertical direction to represent the performance improvement on downstream tasks resulting from domain-adaptive pre-training (see Table 2). Forward Transfer is alternatively referred to as temporal generalization (jin2022lifelong, ) or knowledge transfer (lazaridou2021mind, ) in some literature. In this section, we introduce the evaluation protocols and datasets for continul LLMs.
B.2. Continual LLMs’ Evaluation Protocols
(14) $\Delta R_{t}^{X}$ $\triangleq\tfrac{1}{M}\sum_{m=1}^{M}(R_{t,i}^{X}-b_{i}^{X}).$
B.3. Datasets